The number of chips of different colors in Amy's bag is shown below: 8 blue chips 9 pink chips 1 white chip Amy takes out a chip from the bag randomly without looking. She replaces the chip and then takes out another chip from the bag. What is the probability that Amy takes out a pink chip in both draws?
probability of an event happening = number of ways it can happen
total number of outcomes
Given:
blue chips = 8
pink chips = 9
white chips=1
total = 18
1st draw:
pink chip = 9/18 = 1/2
2nd draw:
pink chip = 9/18 = 1/2
Answer:
81/324
Step-by-step explanation:
9/18 * 9/18 = 81/324